A 163 pounds weight thrombosis patient was ordered to administrate 80 units per kg heparin into a bolus through an infusion. Calculate administered heparin dosage?
Answer #1

Solution: At first units of pounds can be converting into kilograms

163 / 2.2 = 74 kgs

Dose calculation in units: 74 x 80 = 5920units

Calculate mL dosage

1000U : 1ml :: 5920 u : X mL

1000U x XmL = 5920U

X mL = 5920 / 1000 = 5.9 mL bolus

5.9 mL bolus heparin dose was administered